OpenStudy (dtan5457):

if x+y=1 and x^2+y^2=4, what is x^3+y^3

OpenStudy (dtan5457):

weird precalc question i got today no idea how to do it

OpenStudy (dtan5457):

@Nnesha @iambatman

OpenStudy (dtan5457):

@StudyGurl14

OpenStudy (empty):

\[(x+y)^3 = x^3+3x^2y+3xy^2+y^3\] \[(x+y)^2 = x^2+2xy+y^2\] This should be all you need :)

OpenStudy (dtan5457):

im gonna give it try.

OpenStudy (empty):

Sounds good, I'll give you some more hints if you need them later. :D

OpenStudy (dtan5457):

the formula is just the most simplfied versions of x^3+y^3=(x+y)(x^2-xy+y^2) right?

hero (hero):

you have the formula. All you have to do is just input the given Info

OpenStudy (dtan5457):

im just confused on whether im suppose to use the x^3+Y^2 formula or (x+y)^3

hero (hero):

(x+y)(x^2-xy+y^2)=(1)(4-xy)

OpenStudy (empty):

\[(x+y)^3 = x^3+3x^2y+3xy^2+y^3\] \[(x+y)^3 = x^3+y^3 + 3xy(x+y)\] \[(x+y)^3 - 3xy(x+y)= x^3+y^3\] So I arranged it so we have \(x^3+y^3\) which is what we want, and we know that \(x+y=1\) so we plug it in now: \[1^3 - 3xy(1)=1-3xy = x^3+y^3\] What's xy? Let's use this: \[(x+y)^2 = x^2+2xy+y^2\] \[\frac{(x+y)^2 - (x^2+y^2)}{2} = xy\] \[\frac{1^2 - 4}{2} =\frac{-3}{2} = xy\] Now we can plug it in: \[1-3 \frac{-3}{2} = x^3+y^3\]

OpenStudy (dtan5457):

im seeing what you did but if I had started with the formula (x+y)(x^2-xy+y^2) if i plug in from there I get 4-xy? could i continue form here?

OpenStudy (empty):

Yeah It looks like that'd probably work too

OpenStudy (dtan5457):

im not sure what to do from there though

OpenStudy (dtan5457):

do i use the same formula ?

hero (hero):

Use(x+y)^2=x^2+2xy+y^2 from there

hero (hero):

And solve for xy

OpenStudy (dtan5457):

Oh I think I got it. 1^2=4+2xy 1-4=-3 -3/2=2xy xy=-3/2 1(4-3/2)=x^3+y^3

OpenStudy (dtan5457):

can you check for me? and how would you know which formula to use? it seems like you would need more than 1 to compute all the variables

hero (hero):

you will end up with ×y =-3/2 so (1) (4-xy)=4-(-3/2)= 4+3/2= x^3 + y^3

OpenStudy (dtan5457):

oh, dumb negative mistake by me. this is kind of confusing because i never substituted formulas to solve questions like this. normally, how would you know which formulas to use?

OpenStudy (dtan5457):

like how we used x^3+y^3 then to (x+y)^2?

hero (hero):

Try to reason through it using the Known formulas

OpenStudy (dtan5457):

you are right. this is only my first question with this, i just need to practice more. thank you both so much.